OpenStudy (anonymous):

Use Newton’s method to estimate Sqrt(3). Now use linear approximation. Which approach is better?

OpenStudy (freckles):

have you tried one or both methods?

OpenStudy (anonymous):

No, I've been having trouble with Newtons method

OpenStudy (astrophysics):

Newton's method is this right \[\large x_{n+1} = x_n - \frac{ f(x_n) }{ f'(x_n) }\] Damn! The man was connected to the universe!

OpenStudy (anonymous):

Hahaha yes, that's it I guess :D lol

ganeshie8 (ganeshie8):

Finding \(\sqrt{3~}\) is equivalent to solving below equation for positive root \[x^2 = 3\] which is same as finding the roots of below polynomial \[f(x)=x^2-3\tag{1}\]

OpenStudy (freckles):

\[4 \text{ is the closest perfect square to } 3 \\ \text{ and we know } \sqrt{4}=2 \\ x_0=2 \text{ as initial }\]

OpenStudy (freckles):

are you given how many iterations you are to you?

OpenStudy (anonymous):

Right, and nope

OpenStudy (freckles):

because I think we can get closer with this method if we do many

OpenStudy (astrophysics):

I'll give you some info on linear approximations though. So a tangent line to the curve y = f(x) at (a,f(a)) can be used as an approximation of f(x) when x is near a, so a graph will probably demonstrate this better than what I just said lol.  \[y-f(a) = f'(a)(x-a)\] this should look similar. \[y=f'(a)(x-a)+f(a), \implies f(x) \approx f'(a)(x-a)+f(a)\] when x is near a. Therefore we have \[f(x) \approx f'(a) (x-a)+f(a)\] which is the linear approximation! Just testing myself to see if I can derive it also good info for you